OpenStudy (anonymous):

find the limit as x approaches 0 for (sinxcosx)/x

OpenStudy (anonymous):

Use L'Hopital's Rule and notice that the numerator is 1/2(sin(2x)).

OpenStudy (zehanz):

Hi @melisa2695, welcome to OpenStudy! Do you know the standard limit:\[\lim_{x \rightarrow 0}\frac{ \sin x }{ x }=1\] If so, you can write your limit as\[\lim_{x \rightarrow 0}\frac{ \sin x }{ x }\cdot \lim_{x \rightarrow 0}\cos x\]

OpenStudy (zehanz):

If you don't know that standard limit, you could use l'Hopital, as @bfsgd said...

OpenStudy (anonymous):

I try to avoid the standard limits since my memory is appalling......but that would be much easier! @ZeHanz